OpenStudy (anonymous):

Will medal and fan!! How many solutions does the equation 4s + 8 = 3 + 5 + 4s have? Two None One Infinitely many

OpenStudy (mathstudent55):

Have you tried solving it?

OpenStudy (mathstudent55):

To start, add 5 and 3 on the right side. What do you get?

OpenStudy (anonymous):

we know that NONE is wrong, so that one can be eliminated

OpenStudy (anonymous):

if u add both sides u get 12=12, right?

OpenStudy (mathstudent55):

If you end up with no variables and a false statement, there are no solutions. If you end up with no variables and a true statement, there are infinitely many solutions.

OpenStudy (mathstudent55):

\(4s + 8 = \color{red}{3 + 5} + 4s\) Add 3 and 5 together on the right side to get \(4s + 8 = \color{red}{8} + 4s\) Now subtract 8 from both sides: \(4s = 4s\) Divide both sides by 4s: \(1 = 1\) This is a true statement. That means there is an infinite number of solutions. In other words, every real number is a solution of this equation.

OpenStudy (mathstudent55):

@Houdini_Dragon \(4s + 8 = 4s + 8 \ne 12\)

OpenStudy (anonymous):

OOHH-k